# GRASP: A Grid-Based Benchmark for Evaluating Commonsense Spatial Reasoning

### v1

@@ 4 Experiments
The second baseline approach is the greedy search method. This method simulates an agent with perfect knowledge of the grid environment but no planning abilities and no knowledge about the agent constraints except the available movement-related actions. This agent starts from the agent’s current position and uses a Breadth-First Search algorithm to navigate the grid. It considers all four $\mu_{1}$ or eight $\mu_{2}$ possible movement-related actions, depending on the available actions. At each step, it checks the neighboring cells in all directions. The sequence of the neighboring cells in consideration is randomized. If a neighboring cell contains energy, the search stops, and the path to this cell is recorded. If all neighboring cells are either empty or obstacles, the search continues to the next set of neighboring cells. Once the nearest energy cell is found, the recorded path is used to move the agent step-by-step towards that cell. This will continue until a neighboring cell that has energy is too far from the current position, and the remaining number of steps is not enough for the agent to go there and get back to the starting position using the complement of past movement-related actions in reverse order. At this point, the agent simply uses the complement of past movement-related actions in reverse order to get back to the starting cell and drop all the collected energy. The pseudo-code of both methods is provided in Appendix A.2.
Additionally, we evaluate the performance of two LLMs: GPT-3.5-Turbo (i.e., gpt-3.5-turbo-0125) and GPT-4o. All models are used with a temperature setting of 0 to ensure reproducibility. We only test these models with the basic zero-shot prompting method. To construct the prompt used for the LLMs, we utilize all available information, both from the grid environment and agent constraints. Instructions and agent constraints are provided to LLMs in the system prompt, and the grid is provided in the user prompt; the LLMs are required to provide a full list of actions in the response, with no further interaction with the environment. The specific prompts used are detailed in Appendix A.3. The experiments are conducted using one-tenth of the entire GRASP benchmark, for grid instances indexed from 0 to 9. Associated results are reported below.
4.1 Results
$\downarrow$ Control &\&& its Values $\searrow$ Agent $\rightarrow$ Random Walk Greedy Search GPT-3.5- Turbo GPT-4o
Length Energy Length Energy Length Energy Length Energy
Energy Distribution Random 19.00 -1.26 18.75 -0.14 19.70 -1.96 16.60 -1.20
Vertically-skewed 19.00 -1.43 18.65 -0.23 19.59 -1.66 17.40 -1.55
Horizontally-skewed 19.00 -1.33 18.62 -0.07 19.74 -2.12 17.28 -1.43
Cluster 19.00 -1.90 18.81 -0.09 18.58 -2.09 16.65 -1.28
Spiral 19.00 -1.77 18.73 -0.15 19.41 -2.34 17.09 -1.45
Obstacle Yes 19.00 -1.57 18.70 -0.19 19.33 -2.09 17.18 -1.58
No 19.00 -1.51 18.72 -0.09 19.48 -1.98 16.83 -1.18
Starting Position Inner Position 19.00 -1.50 18.73 -0.02 19.38 -1.92 17.04 -1.34
Outer Position 19.00 -1.58 18.70 -0.25 19.43 -2.14 16.96 -1.42
Movement-related Action Set $\mu_{1}$ 19.00 -1.21 18.54 0.80 19.12 -1.89 16.75 -1.13
$\mu_{2}$ 19.00 -1.87 18.89 -1.07 19.69 -2.17 17.26 -1.63
Energy Carrying Limit No Limit 19.00 -1.40 18.73 0.83 19.78 -2.00 16.72 -1.08
2 Units 19.00 -1.68 18.69 -1.10 19.03 -2.06 17.28 -1.68
Energy Cost Per Step 0 Unit 19.00 1.30 18.72 2.66 19.40 1.24 17.15 1.16
0.3 Unit 19.00 -4.38 18.70 -2.93 19.41 -5.30 16.86 -3.92
Average 19.00 -1.54 18.71 -0.14 19.41 -2.03 17.00 -1.38
Table 2 displays the average number of steps taken and units of energy collected and dropped for the two baseline agents (Random Walk and Greedy Search) and the two LLM-based agents (GPT-3.5-Turbo and GPT-4o), evaluated using GRASP. Most Energy results are negative numbers because they are averaged across all control settings, including a 0.3 unit energy cost per step that can lead to a negative energy result if the energy cost exceeds the energy collected. When considering the overall performance of each agent, we find that the Greedy Search agent performs the best across all settings, while GPT-3.5-Turbo performs the worst in 13 out of 14 settings. Additionally, GPT-4o performs better than GPT-3.5-Turbo but worse than the Greedy Search agent and is roughly on par with the Random Walk agent.
Focusing on the performance within each control, we observed Energy differences of 0.4, 0.5, and 0.6 between the two settings of ‘Obstacle,’ ‘Movement-related Action Set,’ and ‘Energy Carrying Limit’ for GPT-4o. This agent performs the best without any obstacles and energy carrying limit while using the action set $\mu_{1}$, which only includes the four regular movements (i.e., UP, DOWN, LEFT, RIGHT) without the diagonal movements. However, these differences are less apparent for GPT-3.5-Turbo, with the maximum difference being only 0.56 units of energy in the ‘Energy Carrying Limit’ control. Notably, GPT-4o achieves, on average, 5.08 units less energy when given the constraint regarding the energy cost per step than when given no such constraint. A similar phenomenon is also observed for the GPT-3.5-Turbo agent. Additionally, the ‘Energy Distribution’ and ‘Starting Position’ controls seem to play a lesser role in affecting the performance of GPT-4o and GPT-3.5-Turbo, as the difference in the amount of energy collected and dropped is less apparent (with the maximum difference being less than 0.35). However, for GPT-3.5-Turbo, the ‘Energy Distribution’ control shows a more significant impact, with a maximum difference of 0.68 between the ‘Spiral’ and ‘Vertically-skewed’ distributions. Turning to the average length of the steps taken by each agent, we observe that GPT-4o takes the least number of steps among the four agents, often between 16 to 17 steps, while GPT-3.5-Turbo takes the most number of steps, taking 19 to 20 steps.
Figure 2: The path formed by the actions of the four agents (Random Walk, Greedy Search, GPT-3.5-Turbo, and GPT-4o) in the shown grid environment. The agents are under the following constraints: unlimited energy carrying limit, zero energy cost per step, and the $\mu_{1}$ movement-related action set (i.e., UP, DOWN, LEFT, RIGHT). Arrows indicate the movement-related actions, and a yellow star next to each cell signifies the agents’ ‘TAKE’ action in that cell.
Figure 2 illustrates a typical action path of the four agents within the shown instance of the grid environment. We observe that the Random Walk agent wanders through the grid without a clear objective and gives the ‘TAKE’ action in cells without energy, while the Greedy Search agent exhibits a more object-oriented behavior, moving directly towards cells with energy and taking it. Focusing on the LLM-based agent, we find that GPT-3.5-Turbo includes cells with energy in its path but often fails to collect energy, repeatedly attempts to move through the cell with an obstacle, and does not return to the starting cell to drop the collected energy. In contrast, the GPT-4o agent demonstrates more intelligent behavior by actively going towards cells with energy and collecting them while avoiding obstacles, yet it still shows inefficient behaviors, such as missing nearby energy cells with energy on them and taking extra and unnecessary steps to achieve the same objectives. These behaviors also reflect the performance of each agent observed in Table 2.
5 Discussion
In this paper, we introduce GRASP, a comprehensive benchmark designed to evaluate the Commonsense Spatial Reasoning (CSR) abilities of LLMs within structured grid environments. GRASP assesses agents’ abilities to navigate the grid environments and collect energy, presenting 16,000 diverse environments characterized by five types of energy distributions, two distinct obstacle configurations, two modes of agent starting position, and three kinds of agent constraints, including two sets of movement-related actions, two limits on the amount of energy that can be carried, and two different energy costs for each step. We evaluate the performance of two classic algorithms (Random Walk and Greedy Search) and advanced LLMs (GPT-3.5-Turbo and GPT-4o) using GRASP. Our findings show that GPT-3.5-Turbo under-performs compared to the Random Walk agent, often unable to successfully navigate the grid, avoid obstacles, and return to the starting point to drop the collected energy, suggesting a lack of CSR abilities. On the other hand, GPT-4o performs similarly to the Random Walk agent but is still less efficient than the Greedy Search agent. While GPT-4o can recognize spatial relationships and navigate around the grid environments, showing a primitive capability of CSR, its use of spatial information to come up with efficient plans remains limited.
Our study comes with several limitations. Firstly, the synthetic nature of GRASP’s grid environments may not fully capture the complexity of real-world CSR tasks, which include more dynamic and unpredictable environments with a greater variety of objects and incomplete or even hidden spatial information. Furthermore, GRASP focuses on commonsense spatial reasoning but does not consider the multi-modal nature of many real-world tasks where abundant visual information is available. Despite the current limitations and potential hallucination problems in LLMs’ multi-modal abilities, a reliable visual interpretation capability could provide the models with a richer set of spatial information than text alone. Additionally, we only consider scenarios with a single agent, whereas real-world environments often involve multiple agents, requiring more sophisticated behaviors, such as interacting and planning together with the other agents. Furthermore, our evaluation was limited to zero-shot prompting methods for LLMs, potentially underestimating their true capabilities.
To address these limitations, GRASP should be expanded to include more dynamic environments that better simulate real-world CSR tasks, preferably developed using real-world data instead of pure synthetic ones. Additionally, with the rapid advancement of multi-modal foundational models, it is crucial to integrate datasets that combine visual and textual spatial information to better capture the complexity of the CSR tasks and provide a more comprehensive evaluation of the models. Extending the benchmark to scenarios involving multiple agents would also more accurately reflect the intricacies of real-world environments. Furthermore, investigating more sophisticated prompting methods could improve the performance of LLMs in these tasks, providing a more holistic evaluation of their commonsense spatial reasoning abilities.
In conclusion, we have presented a novel benchmark for understanding and evaluating the commonsense spatial reasoning abilities of LLMs, albeit with few limitations. Our findings highlight the current limitations of these models in efficiently completing commonsense spatial reasoning tasks and underscore the need for continued research and development. By providing a structured and diverse benchmark, GRASP lays the foundation for future advancements in LLM capabilities, ultimately contributing to the development of more robust and intelligent systems capable of performing a wide range of commonsense spatial reasoning tasks.

### v2

@@ Methods
The greedy search agent starts from the agent’s current position and uses a Breadth-First Search algorithm to navigate the grid. It considers all four $\mu_{1}$ or eight $\mu_{2}$ possible movement-related actions, depending on the available actions. At each step, it checks the neighboring cells in all directions. The sequence of the neighboring cells in consideration is randomized. If a neighboring cell contains energy, the search stops, and the path to this cell is recorded. If all neighboring cells are either empty or obstacles, the search continues to the next set of neighboring cells. Once the nearest energy cell is found, the recorded path is used to move the agent step-by-step toward that cell. This will continue until a neighboring cell that has energy is too far from the current position, and the remaining number of steps is not enough for the agent to go there and get back to the starting position using the complement of past movement-related actions in reverse order. At this point, the agent simply uses the complement of past movement-related actions in reverse order to get back to the starting cell and drop all the collected energy. The pseudo-code of both methods is provided in the Supplementary Material.
Prompts used for evaluating the LLM-based agents
The system prompt:
You are an agent in a grid world. The grid world consists of cells. Each cell may have one unit of energy or no energy at all.{obstacle_prompt} The goal for you is to collect as much energy as possible and put the collected energy back in the cell where you started. You have 20 steps.{movement_prompt} You can collect energy from a cell by being in the cell and TAKE the energy from the cell. If there is no energy in the cell, you cannot take any energy from it.{energy_limit_prompt} You can not move across the boundary of the grid world. You can drop all your energy by DROP.{cost_of_step_prompt} You can use less than 20 steps. Any invalid step will not cause any change in the grid world.
The variables encapsulated in {} are depended upon the specific environment that the agent is in. If the grid has obstacles, then the obstacle_prompt is ‘ Some cells are blocked by obstacles. You cannot move to or through these cells.’ If the grid does not have obstacles, the obstacle_prompt is left as blank. If the agent’s movement-related action set is $\mu_{1}$, the movement_prompt is ‘ For each step, you can choose UP, DOWN, LEFT, RIGHT, TAKE, and DROP. UP allows you to move one cell up in one step. The other movements are similar.’ If the agent’s movement-related action set is $\mu_{2}$, the movement_prompt is ‘ For each step, you can choose UP, DOWN, LEFT, RIGHT, UPLEFT, UPRIGHT, DOWNLEFT, DOWNRIGHT, TAKE, and DROP. UPLEFT allows you to move diagonally one cell up and left in one step. The other movements are similar.’ If the agent has no limit on the number of energy that it can carry, the energy_limit_prompt is left as blank. If the agent has a limit of 2 units of energy that it can carry, the energy_limit_prompt is ‘ You can only carry two unit of energy at a time.’ If the agent has a zero cost per step, then the cost_of_step_prompt is left as blank. If the agent has a cost of 0.3 unit of energy per step, then the cost_of_step_prompt is ‘ Each step costs you 0.3 unit of energy.’
The user prompt:
[…]
The variables encapsulated in are depended upon the specific environment that the agent is in. If the grid has obstacles, then the user_obstacle_prompt is ‘, O is an obstacle’. If the grid does not have obstacles, the user_obstacle_prompt is left as blank. The GRID variable is the text representation of the grid.
If a grid environment has obstacles and the agent is allowed the movement-related actions set $\mu_{1}$ and is restricted to 2 units of energy carrying limit and has a 0.3 unit of energy cost per step, the agent in this grid environment is given the following prompt:
The system message:
You are an agent in a grid world. The grid world consists of cells. Each cell may have one unit of energy or no energy at all. Some cells are blocked by obstacles. You cannot move to or through these cells. The goal for you is to collect as much energy as possible and put the collected energy back in the cell where you started. You have 20 steps. For each step, you can choose UP, DOWN, LEFT, RIGHT, TAKE, and DROP. UP allows you to move one cell up in one step. The other movements are similar. You can collect energy from a cell by being in the cell and TAKE the energy from the cell. If there is no energy in the cell, you cannot take any energy from it. You can only carry two unit of energy at a time. You can not move across the boundary of the grid world. You can drop all your energy by DROP.Each step costs you 0.3 unit of energy. You can use less than 20 steps. Any invalid step will not cause any change in the grid world.
The user message:
[…]
Data Availability
All data generated or analyzed during this study are provided in a GitHub repository accessed at (https://github.com/jasontangzs0/GRASP).
Supplementary Material
Examples of grid instances
Examples of text forms and configurations of the grid environment shown in Figure 2 are reproduced in Figure 3.
Figure 3: (a) The text form of the grid environment expressed in Figure 2. The text form of the grid environment instantiated using: (b) random energy distribution with $p=0.52$, constructed with obstacles and using the inner position as the agent starting position; (c) vertically-skewed energy distribution with $p_{top}=0.38$, constructed without obstacles and using the inner position as the agent starting position; (d) horizontally-skewed energy distribution with $p_{left}=0.37$, constructed with obstacles and using the outer position as the agent starting position; (e) cluster energy distribution with 4 clusters, constructed without obstacles and using the outer position as the agent starting position; (f) spiral energy distribution, constructed with obstacles and using the inner position as the agent starting position.
The pseudo-code for the two baseline agents
Pseudocode of both baseline methods are reproduced under Algorithms 1 and 2 for completeness.
Algorithm 1 Random Walk Method
0:  The set of available movement-related actions $\mu$
1:  Initialize an empty list $actions$
2:  for $i\leftarrow 1$ to $6$ do
3:     Randomly select an action $a$ from $\mu$
4:     Append the selected action $a$ to $actions$
5:     Execute the selected action $a$
6:     Execute TAKE action
7:  end for
8:  for $i\leftarrow 6$ to $1$ step $-1$ do
9:     Execute the complement of $actions[i]$
10:  end for
11:  Execute DROP action
Algorithm 2 Greedy Search Method
0:  The grid environment, agent’s starting position, the set of movement-related actions $\mu$
1:  Initialize an empty list $PastActions$
2:  while The remaining number of step > 0 do
3:     Initialize queue $Q$ with the agent’s starting position
4:     Initialize an empty set $Visited$
5:     Initialize an empty list $Path$
6:     while $Q$ is not empty do
7:        Dequeue current position $current$ from $Q$
8:        Add $current$ to $Visited$
9:        if $current$ contains energy then
10:           Record $Path$ to $current$
11:           Break
12:        end if
13:        Get a list of neighboring cells and their associated actions $(neighbor,action)s$ from $current$ that is one step away using the $action$ from the set of available actions $\mu$
14:        Randomize the order of $(neighbor,action)s$
15:        for each $(neighbor,action)$ in $(neighbor,action)s$ do
16:           if $neighbor$ is not in $Visited$ and $neighbor$ is not an obstacle then
17:              Enqueue $neighbor$ to $Q$
18:              Append $action$ to $Path$
19:           end if
20:        end for
21:     end while
22:     if energy cell found then
23:        if remaining steps are not enough to get to the energy cell, return to the starting position, and drop the energy then
24:           for $i\leftarrow length(PastActions)$ to $1$ step $-1$ do
25:              Execute the complement of $PastActions[i]$
26:           end for
27:           Execute DROP action
28:           Break
29:        else
30:           for $i\leftarrow 1$ to $length(Path)$ do
31:              Execute $Path[i]$
32:              Append $Path[i]$ to $PastActions$.
33:           end for
34:           Execute TAKE action
35:        end if
36:        The remaining number of steps -= length($Path$) + 1
37:     else
38:        for $i\leftarrow length(PastActions)$ to $1$ step $-1$ do
39:           Execute the complement of $PastActions[i]$
40:        end for
41:        Execute DROP action
42:     end if
43:  end while
Selected negative results from GPT-4o-based agent
We provide six different instances where GPT-4o exhibits inefficient behaviors or even fails to obey the rule in Figure 4.
Figure 4: (a) The agent fails to avoid the obstacle. (b) The agent tries to take the energy from a distant cell instead of a nearby cell. (c) The agent is not utilizing all available steps. (d) The agent tries to take the energy from a distant cell instead of a nearby cell, even with an energy cost per step (e) The agent tries to go beyond the boundary of the grid. (f) The agent repeatedly goes back and forth between the same cells.